# Tighter Confidence Bounds for Sequential Kernel Regression

### v1

@@ Appendix A Tail Bound Derivations
Due to Thm. 5.1 of (Flynn et al., 2023), for any $\delta\in(0,1]$, with probability at least $1-\delta$,
@@
We choose $Z_{t}(g_{t})=(g_{t}-f^{*}(x_{t}))\epsilon_{t}$. Using the $\sigma$-sub-Gaussian property of the noise variables, we have
$\psi_{t}(g_{t},\lambda_{t})=\ln\left(\mathbb{E}\left[\mathrm{exp}(\lambda_{t}(g_{t}-f^{*}(x_{t}))\epsilon_{t})|\mathcal{D}_{t-1}\right]\right)\leq\frac{\sigma^{2}\lambda_{t}^{2}(g_{t}-f^{*}(x_{t}))^{2}}{2}.$
@@
$\mathrm{ln}$ $\left(\mathop{\mathbb{E}}_{\boldsymbol{g}_{t}\sim P_{t}}\left[\mathrm{exp}\left(\sum_{s=1}^{t}\lambda(g_{s}-f^{*}(x_{s}))\epsilon_{s}-\frac{\sigma^{2}\lambda_{s}^{2}(g_{s}-f^{*}(x_{s}))^{2}}{2}\right)\right]\right)$ (18)
@@
$(f^{*}(x_{s})-y_{s})^{2}-(g_{s}-y_{s})^{2}$ $=(\epsilon_{s})^{2}-(g_{s}-f^{*}(x_{s})-\epsilon_{s})^{2}$
$=(\epsilon_{s})^{2}-(g_{s}-f^{*}(x_{s}))^{2}+2(g_{s}-f^{*}(x_{s}))\epsilon_{s}-(\epsilon_{s})^{2}$
$=2(g_{s}-f^{*}(x_{s}))\epsilon_{s}-(g_{s}-f^{*}(x_{s}))^{2}.$
@@
$\lambda_{s}(g_{s}-f^{*}(x_{s}))\epsilon_{s}-\frac{\sigma^{2}\lambda_{s}^{2}(g_{s}-f^{*}(x_{s}))^{2}}{2}=\frac{\lambda}{2}(f^{*}(x_{s})-y_{s})^{2}-\frac{\lambda}{2}(g_{s}-y_{s})^{2}+\frac{1}{2}(\lambda_{s}-\sigma^{2}\lambda^{2})(g_{s}-f^{*}(x_{s}))^{2}.$
@@
$\mathrm{ln}\left(\mathop{\mathbb{E}}_{\boldsymbol{g}_{t}\sim P_{t}}\left[\mathrm{exp}\left(\sum_{s=1}^{t}\frac{\lambda}{2}(f^{*}(x_{s})-y_{s})^{2}-\frac{\lambda_{s}}{2}(g_{s}-y_{s})^{2}+\frac{1}{2}(\lambda_{s}-\sigma^{2}\lambda_{s}^{2})(g_{s}-f^{*}(x_{s}))^{2}\right)\right]\right)\leq\ln(1/\delta).$ (19)
@@
$\mathrm{ln}\left(\mathop{\mathbb{E}}_{\boldsymbol{g}_{t}\sim P_{t}}\left[\mathrm{exp}\left(\sum_{s=1}^{t}\frac{1}{2\sigma^{2}}(f^{*}(x_{s})-y_{s})^{2}-\frac{1}{2\sigma^{2}}(g_{s}-y_{s})^{2}\right)\right]\right)\leq\ln(1/\delta).$ (20)
By defining $\boldsymbol{f}_{t}^{*}:=[f^{*}(x_{1}),\dots,f^{*}(x_{t})]^{\top}$ and $\boldsymbol{y}_{t}:=[y_{1},\dots,y_{t}]^{\top}$, rearranging (20), and then writing the resulting inequality using squared norms rather than sums of squares, we obtain
$\|\boldsymbol{f}_{t}^{*}-\boldsymbol{y}_{t}\|_{2}^{2}\leq-2\sigma^{2}\ln\left(\mathop{\mathbb{E}}_{\boldsymbol{g}_{t}\sim P_{t}}\left[\mathrm{exp}\left(-\frac{1}{2\sigma^{2}}\|\boldsymbol{g}_{t}-\boldsymbol{y}_{t}\|_{2}^{2}\right)\right]\right)+2\sigma^{2}\ln(1/\delta).$
@@
Eq. (8) is derived from Eq. (7) using the following Lemma, which is proved in App. B.2 of (Flynn et al., 2023).
Lemma A.1 ((Flynn et al., 2023)).
@@
$\|\boldsymbol{f}_{t}^{*}-\boldsymbol{y}_{t}\|_{2}^{2}$ $\leq-2\sigma^{2}\ln\left(\mathop{\mathbb{E}}_{\boldsymbol{g}_{t}\sim\mathcal{N}(\boldsymbol{0},c\boldsymbol{K}_{t})}\left[\mathrm{exp}\left(-\frac{1}{2\sigma^{2}}\|\boldsymbol{g}_{t}-\boldsymbol{y}_{t}\|_{2}^{2}\right)\right]\right)+2\sigma^{2}\ln(1/\delta).$ (21)
@@ B.1 Useful Lemmas
From $\left\|f^{*}\right\|_{\mathcal{H}}\leq B$ and Eq. (8), for any $\delta\in(0,1]$, with probability at least $1-\delta$, $f^{*}$ satisfies
$\forall t\geq 1,\alpha\geq 0:\quad\|\boldsymbol{f}_{t}^{*}-\boldsymbol{y}_{t}\|_{2}^{2}+\alpha\left\|f^{*}\right\|_{\mathcal{H}}^{2}\leq R_{t}^{2}+\alpha B^{2}.$ (22)
@@
$\left\|\left(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}}\right)^{1/2}(f^{*}-\mu_{\alpha,t})\right\|_{\mathcal{H}}^{2}\leq R_{t}^{2}+\alpha B^{2}-\boldsymbol{y}_{t}^{\top}\boldsymbol{y}_{t}+\boldsymbol{y}_{t}^{\top}\Phi_{t}(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}})^{-1}\Phi_{t}^{\top}\boldsymbol{y}_{t}.$
This means that $f^{*}\in\widetilde{\mathcal{F}}_{t}$ for all $t\geq 1$ and $\alpha>0$. From the definition of the confidence set $\mathcal{F}_{t}$ from Lemma 4.2, every $f\in\mathcal{F}_{t}$ must satisfy $\|\boldsymbol{f}_{t}-\boldsymbol{y}_{t}\|_{2}^{2}+\alpha\left\|f\right\|_{\mathcal{H}}^{2}\leq R_{t}^{2}+\alpha B^{2}$, which means $f\in\widetilde{\mathcal{F}}_{t}$. Therefore, $\mathcal{F}_{t}\subseteq\widetilde{\mathcal{F}}_{t}$. ∎
@@
First, we prove a statement that resembles a representer theorem (Kimeldorf & Wahba, 1971; Schölkopf et al., 2001). We will show that the solution of (23) must be of the form $f(x)=\boldsymbol{k}_{t+1}(x)^{\top}\boldsymbol{w}$, for some weight vector $\boldsymbol{w}\in\mathbb{R}^{t+1}$.
All functions $f\in\mathcal{H}$ can be written in the form $f=f_{\parallel}+f_{\perp}$, where $f_{\parallel}=\boldsymbol{k}_{t+1}(\cdot)^{\top}\boldsymbol{w}$ is in the subspace of $\mathcal{H}$ spanned by $k(\cdot,x_{1}),\dots,k(\cdot,x_{t}),k(\cdot,x)$, and $f_{\perp}$ is orthogonal to this subspace. Since $f_{\perp}$ is orthogonal to the bases $k(\cdot,x_{1}),\dots,k(\cdot,x_{t}),k(\cdot,x)$, we have
@@ B.4 Proof of Theorem 11
Since the Lagrangian $L(f,\eta_{1},\eta_{2})$ is linear in the Lagrange multipliers for every $f\in\mathcal{H}$, the dual function is the maximum over a set of linear functions, which is a convex function (see Eq. (3.7) in Sec. 3.2.3 of (boyd2004convex)). This means that the dual problem is a convex program. We will now show that, for any $\eta_{1}>0$ and $\alpha>0$, there is a closed-form expression for the dual function $g(\eta_{1},\alpha\eta_{1})$. Using Lemma B.1, the Lagrangian evaluated at the Lagrange multipliers $\eta_{1}$ and $\alpha\eta_{1}$ is
@@
B.5 Proof of Corollary 4.6
Proof of Corollary 4.6.
For all $f\in\widetilde{\mathcal{F}}_{t}$ and all $x\in\mathcal{X}$, we have
$|f(x)-\mu_{\alpha,t}(x)|$ $=|k(\cdot,x)^{\top}(f-\mu_{\alpha,t})|$
$=|k(\cdot,x)^{\top}(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}})^{-1/2}(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}})^{1/2}(f-\mu_{\alpha,t})|$
$\leq\sqrt{k(\cdot,x)^{\top}(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}})^{-1}k(\cdot,x)}\left\|(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}})^{1/2}(f-\mu_{\alpha,t})\right\|_{\mathcal{H}}$
$=\frac{1}{\sqrt{\alpha}}\sqrt{k(x,x)-\boldsymbol{k}_{t}(x)^{\top}(\boldsymbol{K}_{t}+\alpha\boldsymbol{I})^{-1}\boldsymbol{k}_{t}(x)}\left\|(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}})^{1/2}(f-\mu_{\alpha,t})\right\|_{\mathcal{H}}$
$\leq\frac{\widetilde{R}_{\alpha,t}}{\sqrt{\alpha}}\sqrt{k(x,x)-\boldsymbol{k}_{t}(x)^{\top}(\boldsymbol{K}_{t}+\alpha\boldsymbol{I})^{-1}\boldsymbol{k}_{t}(x)}.$
The first inequality is due to the Cauchy-Schwarz inequality, and the final equality is due to Lemma B.3. ∎
@@ C.1 Useful Lemmas
Lemma C.2 gives us a simplified expression for $\widetilde{R}_{\alpha,t}$ when $\alpha=\sigma^{2}/c$. This allows us to compare $\widetilde{R}_{\alpha,t}$ the equivalent radius quantities of the confidence bounds in (Abbasi-Yadkori, 2012) and (Chowdhury & Gopalan, 2017).
@@
For any $\delta\in(0,1]$ and any $\lambda>0$, the UCB from Thm. 3.11 of (Abbasi-Yadkori, 2012) states that, with probability at least $1-\delta$,
$\forall t\geq 1:\quad f^{*}(x)\leq\mu_{\lambda,t}(x)+\frac{\widetilde{R}_{\lambda,t}^{\mathrm{AY}}}{\sqrt{\lambda}}\rho_{\lambda,t}(x),$
@@
For any $\delta\in(0,1]$ and any $\eta>0$, the UCB from Thm. 2 of (Chowdhury & Gopalan, 2017) states that, with probability at least $1-\delta$,
$\forall t\geq 1:\quad f^{*}(x)\leq\mu_{1+\eta,t}(x)+\widetilde{R}_{\eta,t}^{\mathrm{IGP}}\rho_{1+\eta,t}(x),$

### v2

@@ Appendix A Tail Bound Derivations
Due to Thm. 5.1 of (Flynn et al.,, 2023), for any $\delta\in(0,1]$, with probability at least $1-\delta$,
@@
We choose $Z_{t}(g_{t})=(g_{t}-f^{\star}(x_{t}))\epsilon_{t}$. Using the $\sigma$-sub-Gaussian property of the noise variables, we have
$\psi_{t}(g_{t},\lambda_{t})=\ln\left(\mathbb{E}\left[\mathrm{exp}(\lambda_{t}(g_{t}-f^{\star}(x_{t}))\epsilon_{t})|\mathcal{D}_{t-1}\right]\right)\leq\frac{\sigma^{2}\lambda_{t}^{2}(g_{t}-f^{\star}(x_{t}))^{2}}{2}.$
@@
$\mathrm{ln}$ $\left(\mathop{\mathbb{E}}_{\boldsymbol{g}_{t}\sim P_{t}}\left[\mathrm{exp}\left(\sum_{s=1}^{t}\lambda(g_{s}-f^{\star}(x_{s}))\epsilon_{s}-\frac{\sigma^{2}\lambda_{s}^{2}(g_{s}-f^{\star}(x_{s}))^{2}}{2}\right)\right]\right)$ (21)
@@
$(f^{\star}(x_{s})-y_{s})^{2}-(g_{s}-y_{s})^{2}$ $=(\epsilon_{s})^{2}-(g_{s}-f^{\star}(x_{s})-\epsilon_{s})^{2}$
$=(\epsilon_{s})^{2}-(g_{s}-f^{\star}(x_{s}))^{2}+2(g_{s}-f^{\star}(x_{s}))\epsilon_{s}-(\epsilon_{s})^{2}$
$=2(g_{s}-f^{\star}(x_{s}))\epsilon_{s}-(g_{s}-f^{\star}(x_{s}))^{2}.$
@@
$\lambda_{s}(g_{s}-f^{\star}(x_{s}))\epsilon_{s}-\frac{\sigma^{2}\lambda_{s}^{2}(g_{s}-f^{\star}(x_{s}))^{2}}{2}=\frac{\lambda}{2}(f^{\star}(x_{s})-y_{s})^{2}-\frac{\lambda}{2}(g_{s}-y_{s})^{2}+\frac{1}{2}(\lambda_{s}-\sigma^{2}\lambda^{2})(g_{s}-f^{\star}(x_{s}))^{2}.$
@@
$\mathrm{ln}\left(\mathop{\mathbb{E}}_{\boldsymbol{g}_{t}\sim P_{t}}\left[\mathrm{exp}\left(\sum_{s=1}^{t}\frac{\lambda}{2}(f^{\star}(x_{s})-y_{s})^{2}-\frac{\lambda_{s}}{2}(g_{s}-y_{s})^{2}+\frac{1}{2}(\lambda_{s}-\sigma^{2}\lambda_{s}^{2})(g_{s}-f^{\star}(x_{s}))^{2}\right)\right]\right)\leq\ln(1/\delta).$ (22)
@@
$\mathrm{ln}\left(\mathop{\mathbb{E}}_{\boldsymbol{g}_{t}\sim P_{t}}\left[\mathrm{exp}\left(\sum_{s=1}^{t}\frac{1}{2\sigma^{2}}(f^{\star}(x_{s})-y_{s})^{2}-\frac{1}{2\sigma^{2}}(g_{s}-y_{s})^{2}\right)\right]\right)\leq\ln(1/\delta).$ (23)
By defining $\boldsymbol{f}_{t}^{\star}:=[f^{\star}(x_{1}),\dots,f^{\star}(x_{t})]^{\top}$ and $\boldsymbol{y}_{t}:=[y_{1},\dots,y_{t}]^{\top}$, rearranging (23), and then writing the resulting inequality using squared norms rather than sums of squares, we obtain
$\|\boldsymbol{f}_{t}^{\star}-\boldsymbol{y}_{t}\|_{2}^{2}\leq-2\sigma^{2}\ln\left(\mathop{\mathbb{E}}_{\boldsymbol{g}_{t}\sim P_{t}}\left[\mathrm{exp}\left(-\frac{1}{2\sigma^{2}}\|\boldsymbol{g}_{t}-\boldsymbol{y}_{t}\|_{2}^{2}\right)\right]\right)+2\sigma^{2}\ln(1/\delta).$
@@
Eq. (8) is derived from Eq. (7) using the following Lemma, which is proved in App. B.2 of (Flynn et al.,, 2023).
Lemma A.1 ((Flynn et al.,, 2023)).
@@
$\|\boldsymbol{f}_{t}^{\star}-\boldsymbol{y}_{t}\|_{2}^{2}$ $\leq-2\sigma^{2}\ln\left(\mathop{\mathbb{E}}_{\boldsymbol{g}_{t}\sim\mathcal{N}(\boldsymbol{0},c\boldsymbol{K}_{t})}\left[\mathrm{exp}\left(-\frac{1}{2\sigma^{2}}\|\boldsymbol{g}_{t}-\boldsymbol{y}_{t}\|_{2}^{2}\right)\right]\right)+2\sigma^{2}\ln(1/\delta).$ (24)
@@ B.1 Useful Lemmas
From $\left\|f^{\star}\right\|_{\mathcal{H}}\leq B$ and Eq. (8), for any $\delta\in(0,1]$, with probability at least $1-\delta$, $f^{\star}$ satisfies
$\forall t\geq 1,\alpha\geq 0:\quad\|\boldsymbol{f}_{t}^{\star}-\boldsymbol{y}_{t}\|_{2}^{2}+\alpha\left\|f^{\star}\right\|_{\mathcal{H}}^{2}\leq R_{t}^{2}+\alpha B^{2}.$ (25)
@@
$\left\|\left(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}}\right)^{1/2}(f^{\star}-\mu_{\alpha,t})\right\|_{\mathcal{H}}^{2}\leq R_{t}^{2}+\alpha B^{2}-\boldsymbol{y}_{t}^{\top}\boldsymbol{y}_{t}+\boldsymbol{y}_{t}^{\top}\Phi_{t}(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}})^{-1}\Phi_{t}^{\top}\boldsymbol{y}_{t}.$
This means that $f^{\star}\in\widetilde{\mathcal{F}}_{t}$ for all $t\geq 1$ and $\alpha>0$. From the definition of the confidence set $\mathcal{F}_{t}$ from Lemma 4.2, every $f\in\mathcal{F}_{t}$ must satisfy $\|\boldsymbol{f}_{t}-\boldsymbol{y}_{t}\|_{2}^{2}+\alpha\left\|f\right\|_{\mathcal{H}}^{2}\leq R_{t}^{2}+\alpha B^{2}$, which means $f\in\widetilde{\mathcal{F}}_{t}$. Therefore, $\mathcal{F}_{t}\subseteq\widetilde{\mathcal{F}}_{t}$. ∎
@@
First, we prove a statement that resembles a representer theorem Kimeldorf and Wahba, (1971); Schölkopf et al., (2001). We will show that the solution of (26) must be of the form $f(x)=\boldsymbol{k}_{t+1}(x)^{\top}\boldsymbol{w}$, for some weight vector $\boldsymbol{w}\in\mathbb{R}^{t+1}$.
All functions $f\in\mathcal{H}$ can be written in the form $f=f_{\parallel}+f_{\perp}$, where $f_{\parallel}=\boldsymbol{k}_{t+1}(\cdot)^{\top}\boldsymbol{w}$ is in the subspace of $\mathcal{H}$ spanned by $k(\cdot,x_{1}),\dots,k(\cdot,x_{t}),k(\cdot,x)$, and $f_{\perp}$ is orthogonal to this subspace. Since $f_{\perp}$ is orthogonal to the basis $k(\cdot,x_{1}),\dots,k(\cdot,x_{t}),k(\cdot,x)$, we have
@@ B.4 Proof of Theorem 11
Since the Lagrangian $L(f,\eta_{1},\eta_{2})$ is linear in the Lagrange multipliers for every $f\in\mathcal{H}$, the dual function is the maximum over a set of linear functions, which is a convex function (see Eq. (3.7) in Sec. 3.2.3 of (Boyd and Vandenberghe,, 2004)). This means that the dual problem is a convex program. We will now show that, for any $\eta_{1}>0$ and $\alpha>0$, there is a closed-form expression for the dual function $g(\eta_{1},\alpha\eta_{1})$. Using Lemma B.1, the Lagrangian evaluated at the Lagrange multipliers $\eta_{1}$ and $\alpha\eta_{1}$ is
@@ C.1 Useful Lemmas
Lemma C.2 gives us a simplified expression for $\widetilde{R}_{\alpha,t}$ when $\alpha=\sigma^{2}/c$. This allows us to compare $\widetilde{R}_{\alpha,t}$ with the equivalent radius quantities of the confidence bounds in (Abbasi-Yadkori,, 2012) and (Chowdhury and Gopalan,, 2017).
@@
For any $\delta\in(0,1]$ and any $\lambda>0$, the UCB from Thm. 3.11 of (Abbasi-Yadkori,, 2012) states that, with probability at least $1-\delta$,
$\forall t\geq 1:\quad f^{\star}(x)\leq\mu_{\lambda,t}(x)+\frac{\widetilde{R}_{\lambda,t}^{\mathrm{AY}}}{\sqrt{\lambda}}\rho_{\lambda,t}(x),$
@@
For any $\delta\in(0,1]$ and any $\eta>0$, the UCB from Thm. 2 of (Chowdhury and Gopalan,, 2017) states that, with probability at least $1-\delta$,
$\forall t\geq 1:\quad f^{\star}(x)\leq\mu_{1+\eta,t}(x)+\widetilde{R}_{\eta,t}^{\mathrm{IGP}}\rho_{1+\eta,t}(x),$
@@
C.3 Why Are Our Confidence Bounds Tighter?
Here, we explain why our analytic confidence bounds in Cor. 4.6 are tighter than the confidence bounds in Thm. 3.11 of (Abbasi-Yadkori,, 2012). The same reasoning explains why our analytic confidence bounds are tighter than those from Thm. 2 of (Chowdhury and Gopalan,, 2017). Both our analytic confidence bounds and the confidence bounds in (Abbasi-Yadkori,, 2012) are derived from a constraint of the form
$\left\|\left(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}}\right)^{1/2}(f^{\star}-\mu_{\alpha,t})\right\|_{\mathcal{H}}\leq R.$
The difference between the confidence bounds is in the radius quantity $R$. The confidence bounds in (Abbasi-Yadkori,, 2012) use a self-normalised concentration inequality (Thm. 3.4 in Abbasi-Yadkori,, 2012), which states that, with high probability,
$\left\|\left(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}}\right)^{-1/2}\Phi_{t}\boldsymbol{\epsilon}_{t}\right\|_{\mathcal{H}}\leq\sigma\sqrt{\ln\left(\det\left(\frac{1}{\alpha}\boldsymbol{K}_{t}+\boldsymbol{I}\right)\right)+2\ln(1/\delta)}.$
Using the definition of $\mu_{\alpha,t}$ and then the triangle inequality, it is shown in (Abbasi-Yadkori,, 2012) that
$\left\|\left(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}}\right)^{1/2}(f^{\star}-\mu_{\alpha,t})\right\|_{\mathcal{H}}$ $=\left\|\left(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}}\right)^{-1/2}(\Phi_{t}\boldsymbol{\epsilon}_{t}-\alpha f^{\star})\right\|_{\mathcal{H}}$
$\leq\left\|\left(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}}\right)^{-1/2}\Phi_{t}\boldsymbol{\epsilon}_{t}\right\|_{\mathcal{H}}+\alpha\left\|\left(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}}\right)^{-1/2}f^{\star}\right\|_{\mathcal{H}}$
$\leq\sigma\sqrt{\ln\left(\det\left(\frac{1}{\alpha}\boldsymbol{K}_{t}+\boldsymbol{I}\right)\right)+2\ln(1/\delta)}+\sqrt{\alpha}B.$
This gives us the expression for the radius in (Abbasi-Yadkori,, 2012). The triangle inequality step causes the log determinant and the RKHS norm to appear under separate square roots. Instead of the self-normalised concentration inequality, we use a bound on $\|\boldsymbol{f}_{t}^{\star}-\boldsymbol{y}_{t}\|_{2}^{2}$ (Eq. (8)), which can be combined with the constraint $\|f^{\star}\|_{\mathcal{H}}\leq B$ more directly via completing the square (Cor. 4.3) or Lagrangian duality.